OpenStudy (korosh23):

Math question! Factor this expression. 6(x2- 4x +4)2 + (x2- 4x + 4) - 1

OpenStudy (korosh23):

Anyone?

myininaya (myininaya):

\[6(x^2-4x+4)^2+(x^2-4x+4)-1\] do you know how to factor: \[6u^2+u-1?\]

OpenStudy (korosh23):

Can you explain how did you do it?

myininaya (myininaya):

I will give you a hint you want to find two factors of a*c that have product a*c and sum b. That is, you want to find two factors of -6 that multiply to be -6 and add up to be 1.

OpenStudy (korosh23):

Exactly. I know how to do it.

OpenStudy (korosh23):

I just do not know how to do the first one.

myininaya (myininaya):

huh?

myininaya (myininaya):

I'm asking you to factor 6u^2+u-1?

myininaya (myininaya):

once you do that we will go back and replace u with (x-2)^2

OpenStudy (korosh23):

Ok.

OpenStudy (triciaal):

@korosh23 use substitution for the expression

OpenStudy (korosh23):

Triciaal what do you mean by substitution?

myininaya (myininaya):

well I think she means subbing x^2-4x+4 with u that is why I'm asking you to factor 6u^2+u-1

OpenStudy (korosh23):

Ya I will factor it in one moment.

OpenStudy (triciaal):

replace instead of (x^2 +4x + 4) write u so the problem became 6u^2 + u -1

OpenStudy (korosh23):

Myininaya the answer is (u+3)(u-2).

myininaya (myininaya):

not exactly

OpenStudy (korosh23):

Triciaal ok I will do it.

myininaya (myininaya):

you found 3(-2)=-6 and 3+(-2)=1 so replace 1u with 3u-2u 6u^2+3u-2u-1 now factor by grouping

OpenStudy (triciaal):

then follow along with what @myininaya is doing

OpenStudy (korosh23):

ok thank you Triciaal. :)

OpenStudy (korosh23):

I learned how to factor by decomposition, criss cross method and bum method, but not group method.

myininaya (myininaya):

haven't heard of any of those methods

myininaya (myininaya):

@triciaal do you those methods

myininaya (myininaya):

know those methods*

OpenStudy (triciaal):

probably by different names

OpenStudy (korosh23):

ya probably.

OpenStudy (triciaal):

example what is PEMDAS? they mean order of operations.

OpenStudy (korosh23):

You have to factor 6u+3u-2u-1 then write (u2-2u)/(3u-6) u(u-2)/3(u-2) (u+3)(u-2) The answer

myininaya (myininaya):

whatever method that is it doesn't work

myininaya (myininaya):

because the answer isn't (u+3)(u-2)

OpenStudy (korosh23):

You asked me to factor 6u+3u-2u-1 and you mentioned that 3u-2u is 1u

myininaya (myininaya):

well not 6u+3u-2u-1 but 6u^2+3u-2u-1 I'm going to show you how to factor by grouping... 6u^2+3u-2u-1 6u^2 and 3u are the first two terms These two terms have 3u in common so 6u^2+3u=3u(2u+1) The last two terms are -2u and -1 these two terms have -1 in cmmon so -2u-1=-1(2u+1) so we have 3u(2u+1)-1(2u+1) Now we have two terms to look at 3u(2u+1) and -1(2u+1) Both of these terms have (2u+1) in common so we can factor out the (2u+1) try to factor the (2u+1) out of 3u(2u+1)-1(2u+1)

OpenStudy (korosh23):

The answer is (3u-1)(2u+1). The method was good.

myininaya (myininaya):

(3u-1)(2u+1) is correct yes factor by grouping does work

myininaya (myininaya):

now replace u with x^2-4x+4

OpenStudy (triciaal):

@korosh23 please learn this method!

myininaya (myininaya):

don't forget to put ( ) around the x^2-4x+4

OpenStudy (korosh23):

Triciaal, for sure.

OpenStudy (korosh23):

ok

OpenStudy (triciaal):

factor by grouping internationally known

OpenStudy (korosh23):

Myininaya, You mean place x2-4x+4 with the u in (3u-1)(2u+1)?

myininaya (myininaya):

\[(3u-1)(2u+1) \\ u=x^2-4x+4 \\ \text{ so replace } u \text{ with } [x^2-4x+4] \\ (3[x^2-4x+4]-1)(2[x^2-4x+4]+1)\]

myininaya (myininaya):

you can simplify the insides of each of the ( )

OpenStudy (korosh23):

ya I am doing the same thing.

OpenStudy (korosh23):

this is not the answer, but am I doing it correct? (3x2-12x+11)(2x2-8x+9)

OpenStudy (triciaal):

before you do that you might want to realize that if (3u-1)and (2u +1) are factors then 3u -1 = 0 or 2u +1 = 0 so that u = 1/3 or u = -1/2 take each of them separate and replace u with the expression and solve for x

OpenStudy (korosh23):

that is a good idea!

OpenStudy (korosh23):

Triciaal is that what Mininaya means ?

OpenStudy (korosh23):

I mean doing the same thing as you mentioned ?

myininaya (myininaya):

I didn't think we were solving for x

myininaya (myininaya):

We were just factoring

OpenStudy (triciaal):

@myininaya

myininaya (myininaya):

@korosh23 's your answer is close but there is a mistake

OpenStudy (korosh23):

what is it Mininaya?

myininaya (myininaya):

no i'm wrong

myininaya (myininaya):

you are right

myininaya (myininaya):

you win @korosh23 !

OpenStudy (korosh23):

I did it correct ?

myininaya (myininaya):

yes it looks great

OpenStudy (korosh23):

Yes the answer is correct. Thank you so much Myininaya and Triciaal. You two helped a lot and I can only give one medal, but I will become fan of you two! :)

myininaya (myininaya):

I'm having cheesecake You can give the medal to @triciaal

OpenStudy (triciaal):

can't eat that thanks

myininaya (myininaya):

lol

OpenStudy (korosh23):

Ok, but both of you deserve medal.

OpenStudy (triciaal):

enjoy fun

OpenStudy (korosh23):

Thanks!